TRIMETHYLAMINE STUDIED BY NON-RIGID GROUP THEORY

\underline{Y G SMEYERS}, Instituto de Estructura de la Materia (CSIC), Serrano 113bis, Madrid, Spain; M VILLA, Departamento de Quimica (CBI, UAM-I). Purisima y Michoacan S/N, Col. Vicentina, Mexico DF. Mexico.

In this work the non-rigid group (NRG) theory has been used to study the complete and detailed structure of trimethylamine restricted NRG. It is shown that the NRG theory is the most suitable one to define properly the 324 symmetry operations contained in the molecular structure. A quite consistent and systematic procedure us used to obtain the structure of the related character table. It has to be pointed out that in the 324 symmetry operations is contained the invertion motion wich by previous Hartree-Fock claculations is showm to have a low energy barrier.